A wire of given material having lenght L and area of cross- section A has a resistance 4ohm. What would be the resistance of anthor wire of same material having lenght l/2 land area of cross section 2A?

Grade:Select Grade

1 Answers

IHate Ppts
6 years ago
We know that the resistance is P(rho) l/a. We can ignore P over here because the material is same in both cases. Now, R=l/a. 4=l/a. In the second case, they are asking resistance if the area is double and length is halved. So, R=l/2 whole divided by 2a.We get R=l/4a. But we l/a=4. So, 4/4=1 and so the answer is 1 ohm.